A motorist is travelling at 15 m/s when he sees a child run into the road. He brakes and
the car comes to rest in 0.75 s.
b) The deceleration of the car is?
1
Expert's answer
2020-05-11T20:10:00-0400

The acceleration of the car

"a=\frac{v_f-v_i}{t}=\frac{0-15}{0.75}=-26.7\:\rm m/s^2"